# Tracking Meets LoRA: Faster Training, Larger Model, Stronger Performance

### v1

@@ 4.2 Quantitative Comparison with the State-of-the-Art Models
In this section, we present a comprehensive evaluation of our proposed LoRAT, benchmarking it against the latest Transformer-based tracking models across five datasets. We follow the official evaluation protocols to ensure fair and accurate comparison. The results are detailed in Table 1. Additionally, we present efficiency comparison with several SOTA Transformer trackers in Table 2.
@@
LaSOT [17] is a large scale benchmark containing 280 test videos. From Table 1, our smallest variant LoRAT-B-224 achieves 0.724 SUC score, already providing a competitive performance. LoRAT-L-224 provides a SUC score of 0.743, beats all other SOTA models. With the current largest backbone network applied in visual tracking, LoRAT-g-384 achieves a breakthrough SUC score of 0.762, largely surpassing the previous SOTA model ARTrack [58] by 3.1%.
[…]
@@
TrackingNet [48] is a large-scale object tracking benchmark. The test split of TrackingNet includes 511 video sequences with first frame annotated. The results in Table 1 demonstrate that all variants of our tracker obtain competitive SUC scores from 0.837 to 0.859. The largest variant LoRAT-g-378 gets the best 0.859 SUC score, outperforming the previous best model by 0.3%.
@@
GOT-10k [31] provides 180 videos for testing. Following the official protocol, we report the results on the models trained only on GOT-10k train split. As reported in Table 1, on the GOT-10k benchmark, the model size greatly impact the final results, more than the input resolution. our LoRAT-g-378 gets a new state-of-the-art AO score of 79.0%.
@@
TNL2k [57] is a recent tracking dataset with 700 test videos. As shown in Table 1, LoRAT-L-378 achieves 0.5% gain over ARTrack [58] at 0.603 SUC score. LoRAT-g-378 achieves the best performance with a SUC score of 0.613.
@@
Efficiency Comparison. We compare LoRAT with other Transformer trackers on efficiency in Table 2, including running speed, floating point operations and number of parameters. Thanks to the highly efficient one-stream architecture, zero inference cost of LoRA, our tracker has a quite impressive running speed to MACs ratio. Our lightest variant LoRAT-B-224, while delivering SOTA performance, operates at an impressive 209 frames per second (fps). The LoRAT-L-378 variant, defeat other trackers on most datasets, but still have a 63 fps. Remarkably, the LoRAT-g-378 variant, leveraging the largest ViT backbone to date, maintains a practical speed of 20 fps. Besides, our tracker also exhibits efficiency in storage requirements. For instance, the visual tracking adaptation module in the LoRAT-g-378 variant, which includes LoRA, token type embedding, and the head, requires only 80 million parameters. This is substantially less than the 1.1 billion parameters of the ViT-g model.
@@ 4.3 Ablation Experiments
[…]
Table 3: Ablation on the performance improvement of LoRA on our tracker. All variants are trained with full parameter fine-tuning. The results are compared with our default settings, i.e. the variants trained with LoRA.
@@
B-224 72.0 ($\downarrow$0.4) 77.5 ($\downarrow$0.4) 48.3 ($\downarrow$0.2) 54.8 ($\downarrow$0.5) 56.8 ($\downarrow$0.5) 58.1 ($\downarrow$0.7)
B-378 72.7 ($\downarrow$0.2) 78.8 ($\downarrow$0.3) 51.3 ($\downarrow$0.4) 58.4 ($\downarrow$0.8) 58.4 ($\downarrow$1.0) 60.3 ($\downarrow$1.6)
L-224 73.5 ($\downarrow$0.8) 79.8 ($\downarrow$1.0) 50.3 ($\downarrow$1.2) 57.3 ($\downarrow$1.4) 59.3 ($\downarrow$0.1) 61.4 ($\downarrow$0.5)
L-378 75.1 ($\downarrow$0.4) 81.8 ($\downarrow$0.4) 53.1 ($\downarrow$0.4) 60.9 ($\downarrow$0.5) 60.4 ($\downarrow$0.4) 63.1 ($\downarrow$0.8)
Performance improvement of LoRA. We compare the performance of LoRAT with full-finetune variants on 4 relatively small variants, as shown in Table 3. LoRA improves the performance of the tracker on all variants, indicating that our model design correctly unleashes the power of parameter-efficient fine-tuning, promoting the performance of visual tracking through mitigating catastrophic forgetting [22] during fine-tuning [14, 63].
Comparison with other PEFT methods. We compare LoRA with other PEFT methods, including a representative of adatper-based PEFT method [50], a representative of prompt tuning-based PEFT method for vision model [32], and a recent zero inference computational cost PEFT method [42]. As shown in Table 4, LoRA outperforms all other PEFT methods on various settings.

### v2

@@ 4.2 Quantitative Comparison with the State-of-the-Art Models
This section comprehensively evaluates the proposed LoRAT, benchmarking it against the latest Transformer-based tracking models across five datasets. The official evaluation protocols are followed to ensure a fair and accurate comparison. The results are detailed in Table 1. Additionally, an efficiency comparison with several state-of-the-art Transformer trackers is presented in Table 2.
@@
LaSOT [18] is a large scale benchmark containing 280 test videos. From Table 1, our smallest variant, LoRAT-B-224, achieves a competitive SUC score of 0.717. LoRAT-L-224 surpasses all other state-of-the-art models with an SUC score of 0.742. Our largest variant, LoRAT-g-384, utilizing the most substantial backbone network ever applied in visual tracking, sets a new benchmark with an SUC score of 0.762, significantly outperforming the previous best, ARTrack [67], by 3.1%.
LaSOText [17] is an extension of LaSOT with 150 additional highly challenging sequences. According to Table 1, LoRAT-B-378, using the ViT-B backbone, achieves a SUC score of 0.531, surpassing the former leader, ARTrack [67], by 0.4%. LoRAT-L-378, with the ViT-L backbone, reaches a new high with an SUC score of 0.566. Our largest variant, LoRAT-g-378, achieves a comparable SUC score of 0.565.
@@
TrackingNet [53] is a large-scale object tracking benchmark. The test split of TrackingNet includes 511 video sequences. The results in Table 1 show that all variants of our tracker achieve competitive SUC scores, ranging from 0.835 to 0.860. Notably, the largest variant, LoRAT-g-378, records the highest SUC score of 0.860, outperforming the previous top-performing model by 0.4%.
@@
GOT-10k [32] provides 180 videos for testing, with a strict protocol that only allows the GOT-10k training split for model training. As reported in Table 1, the size of the model significantly influences performance on this benchmark, much more than the input resolution. our largest variant, LoRAT-g-378, achieves a new state-of-the-art AO score of 78.9%.
@@
TNL2k [66] is a recent tracking dataset with 700 test videos. As Table 1 indicates, LoRAT-L-378 achieves a 2.0% gain over ARTrack [67] with a 0.623 SUC score. LoRAT-g-378 achieves the best performance with an SUC score of 0.627.
@@
Efficiency Comparison. In Table 2, we compare the efficiency of LoRAT against other Transformer-based trackers across key metrics, including running speed, floating point operations, and number of parameters. LoRAT benefits from a highly efficient one-stream architecture and the zero inference cost of LoRA, resulting in an impressive running speed compared to its floating point operations. Our most compact variant, LoRAT-B-224, achieves state-of-the-art (SOTA) performance while operating at 209 frames per second (fps). The intermediate LoRAT-L-378 variant surpasses most competitors, maintaining a speed of 63 fps. Notably, the LoRAT-g-378, equipped with the largest publicly available ViT backbone, still operates at a practical speed of 20 fps. Besides, our tracker also exhibits efficiency in storage requirements. The trainable components of the LoRAT-g-378 variant (which can be considered as a visual tracking adapter for ViT), including LoRA, token type embedding, and the head network, require only 80 million parameters. This represents a significant reduction from the 1.1 billion parameters of the full ViT-g model [54].
@@ 4.3 Ablation Experiments
We conduct detailed ablation studies to dissect the contributions of various components within LoRAT. These models in the experiments are evaluated on three benchmarks: LaSOT [18], LaSOText [17], and TNL2k [66].
Table 3: Ablation on the performance improvement of LoRA-based fine-tuning on our tracker. All variants in the table are trained with full fine-tuning. The results are compared with our default settings, i.e. the variants trained with LoRA.
@@
B-224 70.9 ($\downarrow$0.8) 76.2 ($\downarrow$1.1) 50.0 ($\downarrow$0.3) 57.1 ($\downarrow$0.0) 58.1 ($\downarrow$0.7) 60.1 ($\downarrow$1.2)
B-378 73.2 ($\uparrow$0.3) 78.8 ($\downarrow$0.3) 52.9 ($\downarrow$0.2) 60.5 ($\downarrow$0.1) 60.0 ($\uparrow$0.1) 63.3 ($\downarrow$0.4)
L-224 73.0 ($\downarrow$1.2) 79.3 ($\downarrow$1.6) 52.6 ($\downarrow$0.2) 59.8 ($\downarrow$0.2) 60.8 ($\downarrow$0.3) 64.3 ($\downarrow$0.8)
L-378 74.9 ($\downarrow$0.2) 81.8 ($\downarrow$0.2) 55.2 ($\downarrow$1.4) 63.1 ($\downarrow$2.0) 61.7 ($\downarrow$0.6) 65.7 ($\downarrow$1.3)
Performance improvement of LoRA. We report the performance of LoRAT with full fine-tuning setting in Table 3, with a quantified comparison to LoRA-based fine-tuning. As shown in Table 3, the results demonstrate that LoRA significantly enhances the tracker’s performance across all variants. This performance improvement indicates the effectiveness of our design in leveraging parameter-efficient fine-tuning strategies. By integrating LoRA, we successfully mitigate catastrophic forgetting [23], which is a significant problem when adapting pre-trained models to downstream tasks [15, 74] like visual tracking.
Input embedding. This ablation verifies the effectiveness of the proposed decoupled input embedding scheme. The experiments are conducted on the ViT-L [16] backbone with two input resolution settings. From the results in Table 4, we observe that: 1) Freezing positional embedding shows slightly better performance than the unfreezing one (⑤ vs.  ①). 2) Adopting “shared positional embedding”, “token type embedding” and “foreground indication embedding” (our model) respectively outperforms the “separated positional embedding”, “non-token embedding” and “non-foreground indication embedding” version (⑤ - ⑧). 3) Unfreezing the position embedding will degrade the effectiveness of our proposed foreground indicating embedding (④ vs. ③, ⑧ vs. ⑦) 4) With a higher resolution of input images, foreground indication embedding significantly boost the performance of our tracker (⑩ vs.  ⑨).
Table 4: Ablation on the input embedding configurations. All experiments are conducted on ViT-L [16] backbone with two input resolution settings and with LoRA-based fine-tuning. We examine five factors: 1) Freezing positional embeddings during training; 2) Sharing positional embeddings between template and search region through slicing-based positional embedding adaptation, or two separate positional embeddings are initialized with interpolation-based positional embedding adaptation; 3) Using token type embeddings; 4) Incorporating foreground object indication embeddings; 5) Input resolution (224 vs. 378).
Frozen P. Emb. Shared P. Emb. Type Emb. Foreg. Indic. Res. LaSOT [18] LaSOText [17] TNL2K [66]
SUC P SUC P SUC P
224 73.9 80.4 51.9 59.0 60.6 64.3
224 73.9 80.6 52.7 59.8 60.8 64.8
224 74.2 81.1 52.9 60.2 60.8 64.7
224 74.0 80.8 52.8 60.1 61.2 65.5
224 73.5 80.2 51.8 59.1 60.7 64.4
224 73.8 80.6 53.7 61.4 60.6 64.5
224 74.0 80.7 52.4 59.6 60.7 64.7
224 74.2 80.9 52.8 60.0 61.1 65.1
378 74.4 82.6 55.2 63.2 62.3 67.1
378 75.1 82.0 56.6 65.1 62.3 67.0
MLP-only head network. We evaluate the performance of MLP-only and convolutional head networks under LoRA-based and full fine-tuning settings. As shown in Table 5, the OSTrack [72] convolutional head yields similar competitive results to the MLP-only head in the full fine-tuning setting but fails to converge in the LoRA-based fine-tuning setting. We attribute this to the inductive biases in the convolution layer, which may not satisfy the low "intrinsic dimension" assumption required by low-rank adaptation. Additionally, an MLP-only head offers better computational efficiency.
Table 5: Ablation on head modules. Conv head: Replace the 3-layers MLP-base head in our tracker with 5 stacked Conv-BN-ReLU layers; OSTrack [72] head: Replace the head in our tracker with OSTrack [72] convolution-based head.
[…]
SUC PNorm P SUC PNorm P SUC P fps M
LoRA:
L-224 74.2 83.6 80.9 52.8 64.7 60.0 61.1 65.1 119 336
+ Conv head 4.7 6.4 1.6 4.5 9.4 0.8 0.0 0.1 119 336
+ OSTrack [72] head 5.4 5.1 1.3 4.9 9.0 1.2 0.1 0.1 119 340
Full Fine-tune:
L-224 73.0 81.8 79.3 52.6 64.1 59.8 60.8 64.3 119 307
+ Conv head 71.7 80.5 78.0 51.7 63.1 59.0 58.8 62.0 109 309
+ OSTrack [72] head 72.4 80.9 78.5 51.9 63.2 59.3 59.3 62.6 104 311
Table 6: OSTrack variants implemented using official codes.
Variant Pre-training LaSOT SUC
B-256 MAE 68.7
L-256 MAE 70.3
B-224 DINOv2 66.2
L-224 DINOv2 67.1
Table 7: Training time and max GPU memory usage of our tracker.
Variant Time(h) Memory(GB)
LoRA B-224 5.9 2.4
B-378 12.2 5.7
L-224 10.8 6.6
L-378 32.2 14.9
g-224 22.3 14.0
g-378 60.0 25.8
Full Fine-tuning B-224 5.9 3.2
B-378 12.5 6.5
L-224 12.1 10.0
L-378 34.1 19.1
g-224 29.3 27.1
g-378 Out of Memory
Table 8: Ablation on the impact of ViT backbone with various pre-training methods. The base is B-224 trained with full fine-tuning.
[…]
SUC P SUC P SUC P
MAE [27] 70.3 75.3 49.2 56.0 56.9 58.6
CLIP [57] 70.1 75.0 48.7 55.0 56.8 58.3
EVA-02 [20] 70.7 76.2 49.8 56.9 56.8 58.6
DINOv2 [54] 70.9 76.2 50.0 57.1 58.1 60.1
Table 9: Training time, max training GPU memory usage, max training batch size and inference FPS of our tracker on a single customer-grade NVIDIA RTX 4090 GPU.
Variant Time(h) Memory(GB) Batch Speed (fps)
LoRA B-224 11.0 14.1 128 605
B-378 33.8 20.3 64 285
L-224 34.0 19.5 64 289
L-378 110.3 14.7 16 150
Full Fine-tuning B-224 11.7 16.1 128 605
B-378 36.5 12 32 285
L-224 38.5 13.7 32 289
L-378 117.8 18 16 150
Aligned comparison with OSTrack. We provide an aligned performance comparison with OSTrack [72] by replacing only the backbone of OSTrack. For a fair comparison, OSTrack variants are trained without the early candidate elimination module. Input resolutions of variants are adjusted to align with the patch size of the ViT backbone. Results in Table 6 indicate that all OSTrack variants underperform compared to our model. Besides, OSTrack models using the DINOv2 backbone demonstrate weaker performance compared to their default MAE backbone configuration. These results demonstrate the effectiveness of our model design.
Different Pre-training Methods. We evaluate the performance of four ViT variants, each pre-trained using different methodologies: MAE [27], CLIP [57], EVA02 [20], and DINOv2 [54], when serving as the backbone of our tracker. According to Table 8, the self-supervised DINOv2 pre-training model consistently outperforms other models on all tested datasets. Besides, this superiority, as reflected against the outcomes from Table 6, illustrates that our model design effectively unleashes the power of large-scale pre-training.
Training Efficiency of LoRA. The efficiency of LoRA’s fine-tuning process is evaluated on two different hardware setups: one with 8 NVIDIA V100 GPUs and another with a single NVIDIA 4090 GPU, as shown in Tables 7 and 9, respectively. The results reveal that fine-tuning with LoRA significantly reduces both training time and GPU memory usage, particularly for larger models.
5 Conclusion
In this work, we introduce a novel approach to visual tracking by integrating Low Rank Adaptation (LoRA) within the one-stream tracking framework. Our contributions include two innovative designs that enable effective adaptation of pre-trained models for tracking, significantly reducing the resource requirements. The proposed tracker, LoRAT, not only achieves state-of-the-art performance on multiple benchmarks but also demonstrates the feasibility of training advanced tracking models with manageable resources. This study bridges the gap between advanced large-scale models and resource accessibility, marking a pivotal step towards sustainable advancements in the field of visual tracking.
Acknowledgments. This work is supported by Pengcheng Laboratory Research Project No. PCL2023A08 and No. U20B2052. Heng Fan, Zhipeng Zhang and Haibin Ling receive no financial support for the research, authorship, and/or publication of this article.
Appendix
Appendix 0.A More Implementation Details
This section provides additional details on the implementation of our proposed LoRAT tracker. The hyper-parameters used during training and testing are summarized in Table 10.
0.A.1 General Settings
Input Data. Following common practice in Siamese Trackers [2], the template and search region are cropped from input images. The template has a background area factor of 2, while the search region has a background area factor of 4 for -224 variants and 5 for -378 variants, as in [72].
Training. We apply common data augmentation strategies to the image pairs, including random color jittering and horizontal flipping. Additionally, we employ the 3-Augment [62] strategy, with a minor modification ensuring that the images in a pair are applied with the same transformation. The models are trained using AdamW [49] optimizer, with a learning rate of 1e-4 and a weight decay of 0.1. The learning rate is adjusted with a cosine scheduler. The weight decay for all one-dimensional parameters is set to 0 according to [63]. A 2-epoch linear warmup with an initial learning rate multiplier of 1e-3 is applied to stabilize training. Drop path [36] is applied to -L and -g variants with rates of 0.1 and 0.4, respectively. For full fine-tuning settings, we use layer-wise $lr$ decay values [1] of 0.7/0.8/0.9 for -B/-L/-g variants respectively. The models are trained with automatic mixed precision enabled. The CUDA kernel optimization techniques brought by PyTorch, including torch.compile and memory-efficient attention [56], are also used to reduce GPU memory usage and training time.
Inference. For all performance and efficiency evaluation experiments, automatic mixed precision and memory-efficient attention [56] are enabled.
Table 10: Hyper-parameters used in our models.
Item Value
overall:
template area factor 2
search region area factor 4 (-224) / 5 (-378)
horizontal flip 0.5
color jitter 0.4
batch size 128
epochs 170 / 100 (GOT-10k)
optimizer AdamW
lr 1e-4
weight decay 0.1
drop path 0.0 (B) / 0.1 (L) / 0.4 (g)
clip max norm 1.0
warmup epochs 2
warmup lr mult 1e-3
BCE loss coef 1.0
GIoU loss coef 1.0
hann window panelty 0.45
LoRA:
$r$ 64
full fine-tune:
layer-wise lr decay 0.7 (B) / 0.8 (L) / 0.9 (g)
0.A.2 Head Network
Classification Branch. The classification branch outputs a classification response map $\mathbf{R}\in[0,1]^{N_{S}^{H}\times N_{S}^{W}}$, which assigns a classification score to each point on the feature map. The classification scores at foreground points, those intersecting with the target object, are set to reflect the Intersection over Union (IoU) between the bounding boxes predicted by the regression branch and the ground truth. For background points, the score is targeted to zero. This method of scoring, known as IoU-aware classification score (IACS) as referenced in [73], provides a joint assessment of object presence confidence and localization accuracy.
Bounding Box Regression Branch. Concurrently, the bounding box regression branch generates a map $\mathbf{B}\in[0,1]^{N_{S}^{H}\times N_{S}^{W}\times 4}$, predicting the bounding box coordinates relative to each point on the feature map. At each point $(x_{i},y_{j})$, the branch predicts offsets $(l_{i,j},t_{i,j},r_{i,j},b_{i,j})$, representing the left, top, right, and bottom distances from the point to the edges of the bounding box. The predicted coordinates are calculated as $(x_{i}-l_{i,j},y_{j}-t_{i,j},x_{i}+r_{i,j},y_{j}+b_{i,j})$. The point with the highest classification score is selected as the final bounding box output.
0.A.3 Loss Functions
Classification Loss. We adopt the binary cross-entropy (BCE) loss for classification, formulated as:
$\begin{split}\mathcal{L}_{cls}&=\frac{1}{N_{\text{pos}}}\sum_{i,j}-(\bar{y}_{i,j}\log(\hat{y}_{i,j})+(1-\bar{y}_{i,j})\log(1-\hat{y}_{i,j})),\\ {\textrm{where}}\ \bar{y}&=\left\{\begin{array}[]{ll}\mathrm{IoU}(B_{i,j}^{\text{pred}},B^{\text{gt}})&y_{i,j}=1,\\ 0&y_{i,j}=0,\end{array}\right.\end{split}$ (10)
where $y_{i,j}$ represents the ground-truth label of the location $(i,j)$ on the classification response map $\mathbf{R}$ (1 for foreground, 0 for background), $\hat{y}_{i,j}$ represents the IoU-aware classification score [73], $B_{i,j}^{\text{pred}}$ is the target object bounding box predicted by the bounding box regression branch, and $B^{\text{gt}}$ is the ground-truth bounding box, $N_{\text{pos}}$ is the total number of positive samples on the response map.
Bounding Box Regression Loss. We adopt the GIoU [58] loss for bounding box regression, formulated as:
$\mathcal{L}_{box}=\frac{1}{N_{\text{pos}}}\sum_{i,j}\mathbbm{1}_{{y_{i,j}=1}}\text{GIoU}(B^{\text{gt}},B_{i,j}^{\text{pred}})$ (11)
where $\mathbbm{1}_{{y_{i,j}=1}}$ is the indicator function, ensuring that only positive samples contribute to the loss.
The overall loss function is the sum of the classification loss and the bounding box regression loss, with equal weights (1.0) assigned to both loss terms.
Appendix 0.B Results on Additional Benchmarks
In this section, we report the performance of the tracker on three additional benchmarks: OTB100 [69], NFS [34] and UAV123 [52]. The results are presented in Table 11.
Table 11: Performance of our tracker on three additional benchmarks. Bold indicates the best results.
LoRAT -B-224 LoRAT -B-378 LoRAT -L-224 LoRAT -L-378 LoRAT -g-224 LoRAT -g-378
OTB100 [69] 71.4 71.7 72.3 72.0 72.6 72.6
NFS [34] 64.0 66.6 66.0 66.7 65.0 68.1
UAV123 [52] 72.6 71.9 73.3 72.5 74.2 73.9
Appendix 0.C More Ablation Studies
Table 12: Ablation on the impact of ViT backbone with various pre-training methods.
[…]
SUC P SUC P SUC P
LoRA MAE [27] pre-training:
B-256 69.2 73.9 49.6 56.4 55.9 57.6 215
L-256 72.2 78.8 50.4 57.5 59.7 63.3 125
CLIP [57] pre-training:
B-256 69.2 73.9 49.3 55.4 56.5 58.0 203
L-224 70.3 75.5 50.3 57.1 57.2 59.3 112
EVA-02 [20] pre-training:
B-224 70.4 76.0 49.2 55.7 57.9 60.3 66
L-224 72.6 79.2 52.9 60.7 60.7 64.3 44
DINOv2 [54] pre-training:
B-224 71.7 77.3 50.3 57.1 58.8 61.3 209
L-224 74.2 80.9 52.8 60.0 61.1 65.1 119
Full Fine-tuning MAE [27] pre-training:
B-256 70.3 75.3 49.2 56.0 56.9 58.6 215
L-256 72.9 79.3 50.6 57.6 60.1 63.8 125
CLIP [57] pre-training:
B-256 70.1 75.0 48.7 55.0 56.8 58.3 203
L-224 71.6 76.9 51.2 58.1 57.9 60.0 112
EVA-02 [20] pre-training:
B-224 70.7 76.2 49.8 56.9 56.8 58.6 66
L-224 72.7 79.0 52.2 59.5 58.8 61.7 44
DINOv2 [54] pre-training:
B-224 70.9 76.2 50.0 57.1 58.1 60.1 209
L-224 73.0 79.3 52.6 59.8 60.8 64.3 119
0.C.1 Impact of Different Pre-training Methods
We investigate the impact of pre-training techniques on the performance of our tracker, focusing on four different ViT [16] backbones: MAE [27], CLIP [57], EVA-02 [20] and DINOv2 [54]. A compact version is already provided in Table 8. The resolutions of template and search region are adjusted according to the patch size of the ViT backbone. For models with patch size 14, i.e. DINOv2-B and EVA-02-B, the template size is set to [112 × 112], the search region size is set to [224 × 224]; For models with patch size 16, i.e. MAE-B and CLIP-B, the template and search region sizes are set to [128 × 128] and [256 × 256], respectively.
As detailed in Table 12, LoRA-based fine-tuning exhibits superior performance on larger models (the variants equipped with ViT-L size backbone). Also, LoRA-based fine-tuning works better on the models with stronger pre-training. For instance, the variants pre-trained with DINOv2 work better with LoRA-based fine-tuning across all datasets, underscoring LoRA’s capability to enhance model performance by reducing catastrophic forgetting.
The self-supervised DINOv2 backbone models consistently outperform other variants under both LoRA-based and full fine-tuning configurations. The performance of the variants with EVA-02 pre-training is also competitive, demonstrating the effectiveness of its vision-language pre-training. However, the inference FPS of EVA-02 variants is significantly lower than that of other models, attributed to its several custom structures that differ from the standard ViT architecture.
Table 13: Performance, training time, max training GPU memory usage, inference FPS of various OSTrack variants. All OSTrack variants are implemented using official codes.
[…]
SUC P SUC P SUC P
B-256 MAE 68.7 74.6 46.6 52.4 55.7 56.1 12.6 5.3 124
L-256 MAE 70.3 76.1 46.9 52.7 56.4 57.2 35.0 14.8 52
B-224 DINOv2 66.2 70.7 45.7 50.8 53.5 52.2 11.8 4.8 122
L-224 DINOv2 67.1 71.9 43.5 47.5 54.1 52.9 34.1 13.4 53
0.C.2 Detailed Comparison with OSTrack
We provide detailed performance and efficiency metrics for OSTrack with various additional backbone configurations in Table 13. Compared to our trackers, OSTrack requires significantly longer training times and higher GPU memory usage during training while achieving lower performance.
Table 14: Experiments on LoRA initialization methods. The base model is LoRAT-L-224.
[…]
SUC PNorm P SUC PNorm P SUC P
Kaiming uniform 73.7 83.0 80.2 52.9 64.8 60.2 61.0 65.0
Gaussian normal 73.2 82.4 79.7 52.9 64.9 60.3 60.5 64.4
Truncated normal 74.2 83.6 80.9 52.8 64.7 60.0 61.1 65.1
0.C.3 Performance under Different Initialization Methods of LoRA Weights
Weight initialization is a crucial step for training neural networks. We compare three initialization methods: kaiming uniform (as in the official LoRA implementation), gaussian normal (as in the official paper [31]), and truncated normal distribution (the initialization method used in BERT [14] and ViT [16]). The experiments are conducted on our LoRAT-L-224 variant. From the results in Table 14, we observe that the truncated normal initialization achieves the best overall performance across the three benchmarks.